Circular No. 7477

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000dd IN ANONYMOUS GALAXY A. B. Aazami and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery of an apparent supernova with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Aug. 10.3 (mag ∼ 17.5) and 11.3 UT (mag ∼ 17.4). An earlier image taken on Aug. 3.3 also showed the new object at mag ∼ 18.0. SN 2000dd is located at α = 20^{h}35^{m}00^{s}.10, δ = −23^{o}05′55′′.2 (equinox 2000.0), which is 1′′.6 north of the nucleus of the faint host galaxy, which in turn is 56′′ east and 22′′ north of MCG –4-48-19. A KAIT image of the same field taken on July 27.3 showed nothing at the position of SN 2000dd (limiting mag ∼ 19.0).

NOVAE IN M31 Li also reports the KAIT detection of two apparent novae in M31 =

NGC 224. The first nova, which was discovered by M. Papenkova on an image taken on July 27.5 UT, is located at α = 0^{h}42^{m}44^{s}.0, δ = +41^{o}17′57′′ (equinox 2000.0), which is ∼ 4′′ west and 108′′ north of the nucleus of M31 (the semi-accurate position of the nova is derived from three GSC stars in the field); available magnitudes: July 19.5, > 19.5; 25.5, 19.1; 27.5, 16.9; 28.5, 17.1; Aug. 11.5, 19.0. The second nova, which was dicovered by A. B. Aazami on an image taken on Aug. 10.5, is located at α = 0^{h}42^{m}37^{s}.6, δ = +41^{o}17′37′′ (equinox 2000.0), which is ∼ 75′′ west and 89′′ north of the nucleus of M31; available magnitudes: Aug 6.5, > 19.0; 8.5, 18.6; 10.5, 17.1; 11.5, 17.3. A finding chart of the two novae can be found

at http://astron.berkeley.edu/~bait/2000/n31 2.html.

COMET C/1999 S4 (LINEAR) The production rates given on IAUC 7475 were for observations ob-

tained on July 29.2 UT.

COMET 2P/ENCKE Visual m_{1} estimates: July 29.99 UT, 11.4 (S. E. Shurpakov, Baran,

Belarus, 0.35-m reflector); Aug. 1.45, 11.1 (A. Hale, Cloudcroft, NM, 0.41m reflector); 5.98, 10.8 (V. S. Nevski, Vitebsk, Belarus, 0.30-m reflector); 10.05, 9.7: (M. Reszelski, Szamotuly, Poland, 20×60 binoculars).

2000 August 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green